# Differentially Private Federated Learning: A Systematic Review

### v3

@@ 3.2. LDP-HFL
Building on the optimization of existing LDP mechanisms, many studies utilize parameters shuffling to further mitigate the curse of dimensionality in LDP-HFL. Sun et al. (Sun et al., 2021) proposed the Adaptive-Duchi mechanism, which sets an adaptive perturbation range for each layer of parameters in the model base on Duchi (Duchi et al., 2018). Furthermore, the parameter shuffling method is proposed by them, where they split the model parameters of each client by layer and then shuffle users’ parameters within each layer. If regarding the information in different layers as independent in some scenarios, the privacy budget will no longer need to be split according to the parallel property of DP, further improving utility 131313 Although this method involves the action of shuffling, it does not conform to the concept and definition of the shuffle model. Parameter shuffling only avoids the division of the privacy budget across multiple dimensions under LDP based on the assumption that parameters are independent at each layer, but it does not achieve privacy amplification from LDP to DP.. Zhao et al. (Zhao et al., 2022) also employed parameter shuffling to eliminate the associations between dimensions. In addition, they designed an Adaptive-Harmony mechanism to allocate perturbation interval adaptively to the parameters in each layer of the model. Varun et al. (Varun et al., 2024) proposed SRR-FL, in which they used the Staircase Randomized Response (SRR) mechanism (Wang et al., 2022a) for local perturbations before parameter shuffling. The SRR mechanism differentiates by distributing different perturbation probabilities among various value groups within the domain, with higher perturbation probabilities assigned to values closer to the true value. This approach can enhance the performance and utility of the randomization scheme.
@@
Other research focuses on the heterogeneity of clients in federated scenarios. For the data heterogeneity, Zhang et al. (Zhang et al., 2023b) focused on the performance degradation caused by model heterogeneity in non-i.i.d. data settings, they proposed a personalized federated learning approach (FedBDP) based on Bregman divergence and differential privacy. Their algorithm employs Bregman divergence to quantify the discrepancy between local and global parameters and incorporates it as a regularization term in the update loss function. Additionally, by defining decay coefficients, they dynamically adjust the magnitude of the differential privacy noise for each round. Wang et al. (Wang et al., 2020c) introduced FedLDA, an LDP-based latent Dirichlet allocation (LDA) model tailored for federated learning settings. FedLDA employs an innovative random response mechanism with a prior (RRP), ensuring that the privacy budget remains independent of the dictionary size. This approach enhances accuracy through adaptive and non-uniform sampling processing. In response to the heterogeneous device scenario, Lian et al. (Lian et al., 2022) proposed a browser-based cross-platform federated learning framework called WebFed. To enhance privacy protection, each client adds Laplace noise to the weights of their local models before uploading the training results. And for heterogeneity in privacy budgets, Yang et al. (Yang et al., 2021) proposed PLU-FedOA, Which focus on various privacy requirements of clients. Firstly, they added Laplace noise to parameters locally using different privacy budget $\epsilon_{k}$ to achieve personalized LDP, and then designed a parameters aggregation algorithm for serve to close to the unbiased estimate.
Last but not least, Some methods choose to perturb raw data directly rather than model parameters. Wang et al. (Wang et al., 2022b) treat each data point as an individual user and locally perturb each raw data before uploading it to the edge server. They employ the RAPPOR method (Erlingsson et al., 2014) to encode each feature of the data, followed by individual bit flips. Their experiments demonstrate that this local data perturbation approach effectively withstands data reconstruction attacks while maintaining model efficiency. Unlike directly perturbing the local raw data, Mahawaga et al. (Mahawaga Arachchige et al., 2022) first extract feature vectors from the local datasets using the convolutional and pooling layers in the CNN model. These feature vectors are then converted into flattened 1-D vectors and subjected to unary encoding and perturbation by RAPPOR (Erlingsson et al., 2014) before being uploaded to the server.
@@ 3.3.1. Shuffle model of CL-DP
Specifically, Liu et al. (Liu et al., 2021a) first consider FL in the shuffle model, proposing the FLAME algorithm. In this algorithm, each user samples the dimensions of local updates and perturbs the values in these dimensions for shuffling and aggregation. To simultaneously leverage the privacy amplification benefits of sampling and shuffling, where the privacy amplification from shuffling relies on a certain number of user samples, they further introduce the dummy padding strategy on the shuffler side, ensuring the sampling size remains constant across each dimension. Around the different sampling strategy, Liew et al. (Liew et al., 2022) proposed shuffled check-in protocol in CL-DP, which is similar to self sampling (Girgis et al., 2021a) in the aspect of methodology but achieves $(\epsilon_{l},\delta_{l})$-LDP in theory and deriving a tighter privacy amplification bound with RDP. Beyond the uniform $\epsilon_{l}$-LDP setting for each user, Liu et al. (Liu et al., 2023b) considered heterogeneous privacy budgets that assume different $\epsilon_{k}$-LDP for each client, deriving a tight $(\epsilon_{c},\delta_{c})$-DP bound in the analyzer side based the idea of hiding among clones (Feldman et al., 2022). To improve the utility, they proposed a Clip-Laplace mechanism to bound the Laplace noise in a finite range.
@@ 3.3.2. Shuffle model of SL-DP
In the shuffle model using SL-DP, the neighboring databases for $(\epsilon_{c},\delta_{c})$-DP analysis are defined based on samples rather than clients’ data. To achieve it, a straightforward approach is to perturb each sample’s updates separately, then aggregate and shuffle these samples. Chen et al. (Chen et al., 2023a) adopt this method, further exploring its privacy amplification under a heterogeneous privacy scenario and achieving a tighter privacy bound by $f$-DP (also known as GDP) compared to Liu et al. (Liu et al., 2023b). More typically and beneficially, each user samples only one record to train and reports its corresponding local updates, as shown in Figure 5 (b).
@@ 4.1. Differentially Private VFL
Clients transmit prediction results to the server, posing privacy risks of leaking features (Pasquini et al., 2021). Therefore, some existing work has introduced DP into the model training of VFL to protect features. Chen et al. (Chen et al., 2020) introduced the concept of asynchrony based on VFL. Although there are multiple clients, only one client and server are active at the same time. They introduced noise by “increasing random neuro” and controlled the variance of Gaussian random neurons to satisfy Gaussian differential privacy (GDP). Oh et al. (Oh et al., 2022) replaced traditional neural network models with a vision transformer (ViT) and used split learning (SL) to bypass the large size issue of ViT by shattering data communication at the slicing layer. They also proposed DP-CutMixSL, introducing a Gaussian mechanism to the data smashed by cutout uploaded to the server, making the algorithm satisfy DP. Mao et al. (Mao et al., 2022) proposed a new activation function called ReLU, which transforms private shuffled data and partial losses into random responses in forward and backward propagation to prevent attribute inference and data reconstruction in SplitNN. Based on this, they introduced differential privacy and added Laplace noise to the forward propagation results in a random response (RR) manner, achieving a fine-grained privacy budget allocation scheme for SplitNN. Tian et al. (Tian et al., 2023) introduced ”bucket” into the scenario of combining VFL with gradient boosting decision tree (GBDT), allowing the server to only know which bucket a sample belongs to, without knowing the order of the sample in the buckets. By moving samples between buckets with a certain probability and adding RR perturbation to each bucket, the samples are protected by DP. Li et al. (Li et al., 2022a) proposed condensed-LDP (CLDP) in the context of VFL combined with tree boosting and designed three order-preserving desensitization algorithms for feature information to achieve privacy protection.
@@
(b) A example of differentially private HFL.
Figure 6. Differentially private VFL and HFL.
Table 3. An overview study of Differentially Private FL.
Federated
Scenario
Publications
DP
Model
Neighborhood
Level
Perturbation
Mechanism
Composition
Mechanism1
Downsteam
Tasks2
Model
Archiecture
Clients
Number
$\epsilon$ $\delta$
Huang et al. (Huang et al., 2019) Gaussian AC Regression LR - [0.01,0.2] $[10^{-3},10^{-6}]$
Huang et al. (Huang et al., 2020) Gaussian, Laplace AC Classification Shallow CNN 10,100,1000 [0.2,8] $[10^{-2},10^{-5}]$
Wei et al. (Wei et al., 2020) Gaussian MA Classification Shallow CNN,LSTM [10,20] [0.12,2] $[10^{-2},10^{-5}]$
Wei et al. (Wei et al., 2021) Gaussian MA Classification Shallow CNN 50 [4,20] $10^{-3}$
Liu et al. (Liu et al., 2021b) Gaussian GDP Classification Shallow CNN 100 [10,100] $10^{-3}$
Zheng et al. (Zheng et al., 2021) Gaussian GDP Classification Shallow CNN 100 [10,100] $10^{-3}$
Noble et al. (Noble et al., 2022) Gaussian RDP Classification Shallow CNN 10 [3,13] $10^{-6}$
Fu et al. (Fu et al., 2022a) Gaussian RDP Classification Shallow CNN 10 [2,6] $10^{-5}$
Li et al. (Li et al., 2022c) Gaussian MA Classification LR, Shallow CNN 10 [1,16] $10^{-3}$
Ryu et al. (Ryu and Kim, 2022) Gaussian AC Classification LR [10,195] [0.05,5] $10^{-6}$
Ling et al. (Ling et al., 2023) Gaussian RDP Classification Shallow CNN 10 [1.5,5.5] $10^{-5}$
Xiang et al. (Xiang et al., 2023) Gaussian MA Classification Shallow CNN,LSTM [10,20] [0.12,2] $[10^{-2},10^{-5}]$
Ruan et al. (Ruan et al., 2023) DP SL Gaussian RDP Classification Shallow CNN, LSTM [3,10] [0.25,2] $[10^{-4},10^{-5}]$
Geyer et al. (Geyer et al., 2017) Gaussian MA Classification Shallow CNN 100, 1000, 10000 8 $[10^{-3},10^{-6}]$
Mamahan et al. (McMahan et al., 2017b) Gaussian MA Classification LSTM [100,763430] [2.0,4.6] $10^{-9}$
Andrew et al. (Andrew et al., 2021) Gaussian RDP Classification Shallow CNN [500,342000] [0.035,5] [$\frac{1}{500}$,$\frac{1}{342000}$]
Zhang et al. (Zhang et al., 2022) Gaussian MA Classification Shallow CNN, ResNet-18 1920 [1.5,5] $10^{-5}$
Cheng et al. (Cheng et al., 2022) Gaussian MA Classification Shallow CNN, ResNet-18 3400 [2,8] $\frac{1}{3400}$
Bietti et al. (Bietti et al., 2022) Gaussian MA Classification Shallow CNN 1000 [0.1,1000] $10^{-4}$
Shi et al. (Shi et al., 2023) Gaussian RDP Classification ResNet-18 500 [4,10] $\frac{1}{500}$
Xu et al. (Xu et al., 2023a) Gaussian RDP Classification ResNet-50 [1262,9896000] [10,20] $10^{-7}$
Yang et al. (Yang et al., 2023b) CL Gaussian RDP Classification Shallow CNN 50 [2,16] $10^{-3}$
Agarwal et al. (Agarwal et al., 2018) Binomial AC Classification LR 25M [2,4] $10^{-9}$
Wang et al. (Wang et al., 2020a) Discrete Gaussian RDP Classification Shallow CNN 100K [2,4] $10^{-5}$
Kerkouche et al. (Kerkouche et al., 2021) Gaussian MA Classification Shallow CNN [5011,6000] [0.5,1] $10^{-5}$
Kairouz et al. (Kairouz et al., 2021a) Discrete Gaussian zCDP Classification Shallow CNN 3400 [3,10] $\frac{1}{3400}$
Agarwal et al. (Agarwal et al., 2021) Skellam RDP Classification Shallow CNN 1000k [5,20] $10^{-6}$
Chen et al. (Chen et al., 2022b) Poisson Binomial RDP Classification LR 1000 [0.5,6] $10^{-5}$
Stevens et al. (Stevens et al., 2022) LWE RDP Classification Shallow CNN [500,1000] [2,8] $10^{-5}$
Chen et al. (Chen et al., 2022a) CL with SA Discrete Gaussian RDP Classification Shallow CNN [100,1000] [0,10] $10^{-2}$
Naseri et al. (Naseri et al., 2020) SL, CL Gaussian RDP Classification Shallow CNN,LSTM [100,660120] [1.2,10.7] $10^{-5}$
Yang et al. (Yang et al., 2023c) DP SL, CL, CL with SA Gaussian, Skellam RDP Classification Shallow CNN [40,500] [2,8] $10^{-3}$
Triastcyn et al. (Triastcyn and Faltings, 2019) Bayesian DP SL, CL Gaussian RDP Classification ResNet-50 [100,10000] [0.2,4] $[10^{-3},10^{-6}]$
Wang et al. (Wang et al., 2019b) PM BC Classification LR, SVM 4M [0.5,4] 0
Liu et al. (Liu et al., 2020a) RR, PM BC Classification LR, SVM 4W-10W [0.5,16] 0
Zhao et al. (Zhao et al., 2020b) Three output, PM-SUB BC Classification LR, SVM 4M [0.5,4] 0
Wang et al. (Wang et al., 2020c) RRP AC Topic Modeling LDA 150 [5,8] [0.05,0.5]
Yang et al. (Yang et al., 2021) Laplace BC Classification Shallow CNN [200,1000] [1,10] 0
Sun et al. (Sun et al., 2021) Adaptive-Duchi BC Classification Shallow CNN [100,500] [1,5] 0
Lian et al. (Lian et al., 2022) Laplace BC Classification Shallow CNN 5 [3,6] 0
Mahawaga et al. (Mahawaga Arachchige et al., 2022) RAPPOR BC Classification Shallow CNN [2,100] [0.5,10] 0
Wang et al. (Wang et al., 2022b) RAPPOR BC Classification LR [500,1800] [0.1,10] 0
Jiang et al. (Jiang et al., 2022) EM BC Classification Shallow CNN [100,750] [0.5,12] 0
Li et al. (Li et al., 2022b) Laplace BC Classification Shallow CNN 100 78.5 0
Zhao et al. (Zhao et al., 2022) Adaptive-Harmony BC Classification Shallow CNN 200 [1,10] 0
Wang et al. (Wang et al., 2023a) EM, DMP-UE BC Classification Shallow CNN [10,50] [0.1,1] 0
Zhang et al. (Zhang et al., 2023b) Gaussian AC Classification LR, Shallow CNN 100 [3,30] -
Varun et al. (Varun et al., 2024) LDP - SRR BC Classification Shallow CNN 100 [1,10] 0
Truex et al. (Truex et al., 2020) Condensed LDP - EM BC Classification Shallow CNN 50 1 0
Liu et al. (Liu et al., 2021a) Laplace,Shuffle BC, AC Classification LR 1000 4.696 $5*10^{-6}$
Liew et al. (Liew et al., 2022) Harmony,Shuffle RDP Classification Shallow CNN [50000,60000] [2.8]
Liu et al. (Liu et al., 2023b) CL Clipped-Laplace,Shuffle AC Classification LR 10000 25.6 $10^{-8}$
Girgis et al. (Girgis et al., 2021b) Laplace,Shuffle AC Classification Shallow CNN 60000 [1,10] $10^{-5}$
Horizontal Chen et al. (Chen et al., 2023a) Shuffle Model SL Duchi,Shuffle GDP Classification Shallow CNN 100 [0.5, 100] $10^{-5}$
Oh et al. (Oh et al., 2022) DP Gaussian RDP Classification VGG-16 10 [1,40] -
Yang et al. (Yang et al., 2022) Laplace, KRR BC Classification Shallow CNN 2 1 0
Takahashi et al. (Takahashi et al., 2023) Label DP SL KRR BC Classification GBDT 3 [0.1,2.0] -
Chen et al. (Chen et al., 2020) Gaussian GDP Classification Shallow CNN [3,8] - -
Wang et al. (Wang et al., 2020b) Gaussian AC Classification Shallow CNN 2 [0.001, 10] $10^{-2}$
Wu et al. (Wu et al., 2020) DP CL Laplace BC Classification GBDT [2,10] - -
Tian et al. (tian2020federboost) RR BC Classification GBDT 3 4 0
Mao et al. (Mao et al., 2022) LDP - Laplace, RR BC Classification Shallow CNN 5 [0.1,4.0] 0
Vertical Li et al. (Li et al., 2022a) Condensed LDP - Discrete Laplace BC Classification GBDT 2 [0.64,2.56] 0
Papernot er al. (Papernot et al., 2017) Laplace MA Classification Shallow CNN 2 [2.04,8.19] $[10^{-5},10^{-6}]$
Papernot er al. (Papernot et al., 2018) Gaussian RDP Classification Resnet-18 2 [0.59,8.03] $10^{-8}$
Sun et al. (Sun and Lyu, 2020) Random Sampling AC Classification Shallow CNN 6 [0.003,0.65] [0.006,0.65]
Tian et al. (Tian et al., 2022) Gaussian GDP Text Generation GPT-2 2000 [3,5] $10^{-6}$
Hoech et al. (Hoech et al., 2022) Gaussian AC Classification Resnet-18 20 [0.1,0.5]
Wan et al. (Wan et al., 2023) SL Gaussian AC Recommendection DeepFM 2 [0.05, 10] -
Pan et al. (Pan et al., 2021a) Gaussian RDP Classification Resnet-18 100 [0.95,9.03]
Dodwadmath et al. (Dodwadmath and Stich, 2022) DP CL Laplace MA Classification Shallow CNN 10 [11.75,20] $10^{-5}$
Transform Qi et al. (Qi et al., 2023) LDP - KRR BC Classification Shallow CNN [2,5] [2,7] 0
  • 1
    1. BC=Basic Sequential Composition Theory, AC=Advanced Sequential Composition Theory.
  • 2

### v4

@@ 3.2. LDP-HFL
Building on the optimization of existing LDP mechanisms, many studies utilize parameters shuffling to further mitigate the curse of dimensionality in LDP-HFL. Sun et al. (Sun et al., 2021) proposed the Adaptive-Duchi mechanism, which sets an adaptive perturbation range for each layer of parameters in the model base on Duchi (Duchi et al., 2018). Furthermore, the parameter shuffling method is proposed by them, where they split the model parameters of each client by layer and then shuffle users’ parameters within each layer. If regarding the information in different layers as independent in some scenarios, the privacy budget will no longer need to be split according to the parallel property of DP, further improving utility  [^8] Although this method involves the action of shuffling, it does not conform to the concept and definition of the shuffle model. Parameter shuffling only avoids the division of the privacy budget across multiple dimensions under LDP based on the assumption that parameters are independent at each layer, but it does not achieve privacy amplification from LDP to DP.. Zhao et al. (Zhao et al., 2022) also employed parameter shuffling to eliminate the associations between dimensions. In addition, they designed an Adaptive-Harmony mechanism to allocate perturbation interval adaptively to the parameters in each layer of the model. Varun et al. (Varun et al., 2024) proposed SRR-FL, in which they used the Staircase Randomized Response (SRR) mechanism (Wang et al., 2022a) for local perturbations before parameter shuffling. The SRR mechanism differentiates by distributing different perturbation probabilities among various value groups within the domain, with higher perturbation probabilities assigned to values closer to the true value. This approach can enhance the performance and utility of the randomization scheme.
@@
Other research focuses on the heterogeneity of clients in federated scenarios. For the data heterogeneity, Zhang et al. (Zhang et al., 2023b) focused on the performance degradation caused by model heterogeneity in non-i.i.d. data settings, they proposed a personalized federated learning approach (FedBDP) based on Bregman divergence and differential privacy. Their algorithm employs Bregman divergence to quantify the discrepancy between local and global parameters and incorporates it as a regularization term in the update loss function. Additionally, by defining decay coefficients, they dynamically adjust the magnitude of the differential privacy noise for each round. Wang et al. (Wang et al., 2020c) introduced FedLDA, an LDP-based latent Dirichlet allocation (LDA) model tailored for federated learning settings. FedLDA employs an innovative random response mechanism with a prior (RRP), ensuring that the privacy budget remains independent of the dictionary size. This approach enhances accuracy through adaptive and non-uniform sampling processing. In response to the heterogeneous device scenario, Lian et al. (Lian et al., 2022) proposed a browser-based cross-platform federated learning framework called WebFed. To enhance privacy protection, each client adds Laplace noise to the weights of their local models before uploading the training results. And for heterogeneity in privacy budgets, Yang et al. (Yang et al., 2021) proposed PLU-FedOA, Which focus on various privacy requirements of clients. Firstly, they added Laplace noise to parameters locally using different privacy budget $\epsilon_{k}$ to achieve personalized LDP, and then designed a parameters aggregation algorithm for serve to close to the unbiased estimate. Zhang et al. (Zhang et al., 2024) improved model performance by dynamically allocating the privacy budget at each federated aggregation round.
Last but not least, Some methods choose to perturb raw data directly rather than model parameters. Wang et al. (Wang et al., 2022b) treat each data point as an individual user and locally perturb each raw data before uploading it to the edge server. They employ the RAPPOR method (Erlingsson et al., 2014) to encode each feature of the data, followed by individual bit flips. Their experiments demonstrate that this local data perturbation approach effectively withstands data reconstruction attacks while maintaining model efficiency. Unlike directly perturbing the local raw data, Mahawaga et al. (Mahawaga Arachchige et al., 2022) first extract feature vectors from the local datasets using the convolutional and pooling layers in the CNN model. These feature vectors are then converted into flattened 1-D vectors and subjected to unary encoding and perturbation by RAPPOR (Erlingsson et al., 2014) before being uploaded to the server. Wang et al. (Wang and Xu, 2019; Wang et al., 2023b) analyse the theoretical error in generalized linear models under LDP.
@@ 3.3.1. Shuffle model of CL-DP
Specifically, Liu et al. (Liu et al., 2021a) first consider FL in the shuffle model, proposing the FLAME algorithm. In this algorithm, each user samples the dimensions of local updates and perturbs the values in these dimensions for shuffling and aggregation. To simultaneously leverage the privacy amplification benefits of sampling and shuffling, where the privacy amplification from shuffling relies on a certain number of user samples, they further introduce the dummy padding strategy on the shuffler side, ensuring the sampling size remains constant across each dimension. Around the different sampling strategy, Liew et al. (Liew et al., 2022) proposed shuffled check-in protocol in CL-DP, which is similar to self sampling (Girgis et al., 2021a) in the aspect of methodology but achieves $(\epsilon_{l},\delta_{l})$-LDP in theory and deriving a tighter privacy amplification bound with RDP. Beyond the uniform $\epsilon_{l}$-LDP setting for each user, Liu et al. (Liu et al., 2023) considered heterogeneous privacy budgets that assume different $\epsilon_{k}$-LDP for each client, deriving a tight $(\epsilon_{c},\delta_{c})$-DP bound in the analyzer side based the idea of hiding among clones (Feldman et al., 2022). To improve the utility, they proposed a Clip-Laplace mechanism to bound the Laplace noise in a finite range.
@@ 3.3.2. Shuffle model of SL-DP
In the shuffle model using SL-DP, the neighboring databases for $(\epsilon_{c},\delta_{c})$-DP analysis are defined based on samples rather than clients’ data. To achieve it, a straightforward approach is to perturb each sample’s updates separately, then aggregate and shuffle these samples. Chen et al. (Chen et al., 2023a) adopt this method, further exploring its privacy amplification under a heterogeneous privacy scenario and achieving a tighter privacy bound by $f$-DP (also known as GDP) compared to Liu et al. (Liu et al., 2023). More typically and beneficially, each user samples only one record to train and reports its corresponding local updates, as shown in Figure 5 (b).
@@ 4.1. Differentially Private VFL
Clients transmit prediction results to the server, posing privacy risks of leaking features (Pasquini et al., 2021). Therefore, some existing work has introduced DP into the model training of VFL to protect features. Chen et al. (Chen et al., 2020) introduced the concept of asynchrony based on VFL. Although there are multiple clients, only one client and server are active at the same time. They introduced noise by “increasing random neuro” and controlled the variance of Gaussian random neurons to satisfy Gaussian differential privacy (GDP). Oh et al. (Oh et al., 2022) replaced traditional neural network models with a vision transformer (ViT) and used split learning (SL) to bypass the large size issue of ViT by shattering data communication at the slicing layer. They also proposed DP-CutMixSL, introducing a Gaussian mechanism to the data smashed by cutout uploaded to the server, making the algorithm satisfy DP. Mao et al. (Mao et al., 2022) proposed a new activation function called ReLU, which transforms private shuffled data and partial losses into random responses in forward and backward propagation to prevent attribute inference and data reconstruction in SplitNN. Based on this, they introduced differential privacy and added Laplace noise to the forward propagation results in a random response (RR) manner, achieving a fine-grained privacy budget allocation scheme for SplitNN. Tian et al. (Tian et al., 2023) introduced ”bucket” into the scenario of combining VFL with gradient boosting decision tree (GBDT), allowing the server to only know which bucket a sample belongs to, without knowing the order of the sample in the buckets. By moving samples between buckets with a certain probability and adding RR perturbation to each bucket, the samples are protected by DP. Li et al. (Li et al., 2022a) proposed condensed-LDP (CLDP) in the context of VFL combined with tree boosting and designed three order-preserving desensitization algorithms for feature information to achieve privacy protection. Liao et al. (Liao et al., 2025) proposed MFVFL, a tensor decomposition-based vertical federated learning framework that resolves feature missingness through client learnable imputation and server low-rank decomposition, while its privacy-preserving variant MFVFL-DP employs tensor robust PCA to mitigate differential privacy noise impact.
@@
(b) A example of differentially private TFL.
Figure 6. Differentially private VFL and TFL.